### v1

@@ III-C Pulse Optimization
To maximize the number of covert bits that can be reliably conveyed, we jointly tune four pulse-design parameters: the pilot-pulse length $n^{(p)}_{s}$, data-pulse length $n^{(q)}_{s}$, pilot-pulse magnitude $c_{p}$, and data-pulse magnitude $c_{q}$. We model the relationship between these parameters and the attainable covert bit count as a black-box objective, apply Bayesian optimization to navigate the parameter space efficiently, and use our hardware testbed to measure the covert bits delivered for each candidate configuration.
Bayesian optimization iteratively explores the four-dimensional design space $\vec{d}=\left(n^{(p)}_{s},n^{(q)}_{s},c_{p},c_{q}\right)$ by maintaining a surrogate model of the unknown mapping from $\vec{d}$ to attainable covert bit count. We leverage Ax library in python [48] to perform this optimization. The three steps that follow are the standard Bayesian-optimization loop [49, Ch. 31]; we restate them here.
@@
  • Measure : using $\vec{d}_{u}$, the radio transmits a modified version of the data packet (Section III-D): every fifth pulse slot in the Alice-on segment carries a random bit and the Alice-off segment is removed. Number of attainable covert bits is recorded over $100$ one-second trials. The average number of attainable covert bits for a given parameter set is $r=2\,C_{\mathrm{bsc}}\,n_{p}\,\alpha_{n}$ bits over $n$ uses of the channel. In this experiment, $n$ corresponds to the number of samples available in a one second transmission. This experiment yields a sample mean $\bar{r}_{u}$ and a standard error $s_{\bar{r}_{u}}$.
@@
After 200 iterations we report $\vec{d}^{\star}=\operatorname*{arg\,max}_{1\leq l\leq 200}\bar{r}_{l},$ the configuration that achieved the highest observed mean throughput.
We use a Gaussian-process surrogate – commonly adopted in Bayesian optimization – and briefly restate only its essentials here. For a set of $l$ different input parameter configurations $\vec{d}_{1:l}=[\vec{d}_{1},\vec{d}_{2},...\vec{d}_{l}]$, the model predicts $l$ corresponding covert throughput values sampled from the distribution $\mathcal{GP}\left(0,\,k(\vec{d}_{1:l},\vec{d}_{1:l})\right)$ [50], where $k$ is a covariance function, which in our case, is set to the Matérn-5/2 function [50, Ch. 4]. This model can also be used to predict the throughput given priors, i.e. given u1u-1 pairs of $(\vec{d},\bar{r},s_{\bar{r}})$, the model can predict the throughput $\hat{r}_{u}$ for some new parameter set $\vec{d}_{u}$. This model prediction is distributed as $\hat{r}_{u}\mid\bar{r}_{1:(u-1)},\,\vec{d}_{1:u},\,s_{\bar{r}_{1:(u-1)}}\sim\mathcal{N}(\mu_{u},\sigma_{u}^{2})$, where $\mu_{u}$ and $\sigma_{u}$ can be directly calculated. At iteration $u$ of the Bayesian optimization, configuration $d_{u}$ that maximizes expected improvement is proposed for measurement. Expected improvement for a candidate parameter configuration $d_{u}^{\prime}$ is defined as $\mathbb{E}[\hat{r}_{u}^{\prime}\mid\hat{r}_{u}^{\prime}>\bar{r}^{\star}_{1:(u-1)}]$, where $\hat{r}^{\star}_{1:(u-1)}$ is the highest measured throughput for the first u1u-1 trials, i.e. $r^{\star}_{1:(u-1)}=\max_{1\leq v\leq(u-1)}\bar{r}_{v}$
The resulting optimum $\left(n^{(p)}_{s},n^{(q)}_{s},c_{p},c_{q}\right)$ is adopted for all experiments reported henceforth.
@@ III-D Transmission Structure for Experimental Trials
In each experimental trial, Alice transmits the three-segment packet shown in Fig. 5: a non-covert preamble, an Alice-on segment, and an Alice-off (noise-only) segment. An additive-white-Gaussian-noise (AWGN) generator operates continuously throughout the entire experiment. The Alice-on and Alice-off segments each last $T$ seconds, while the preamble marks the start of every trial and facilitates synchronization at Bob and Willie.
The preamble consists of the 13-bit Barker sequence repeated five times (65 bits in total). Each bit is BPSK-modulated and occupies 200 samples; the first sample carrying the modulated symbol followed by 199 zero-padding samples. The resulting 13 000-sample sequence is then pulse-shaped with a root-raised-cosine (RRC) filter with roll-off factor $\beta=0.35$ spanning 12 symbols, represented digitally by $12\times 200+1=2401$ taps. After filtering, the preamble length is $65\times 200+(2401-1)=15,400$ samples.
@@ III-E Experiment design
We select eight linearly-spaced values of Alice’s transmission duration $T\in[1,8]$ s. For each value of $T$, we conduct $N=150$ independent trials using distinct transmission packets described in Section III-D. Before we begin our experiment, we estimate Willie’s SNR using a $500$ 0.5 second calibration transmission with a modified version of the data-packet described in subsection III-D: every fifth pulse slot in the Alice-on segment is used to transmit a random bit and Alice-off segment is deleted. The SNR is computed from the mean power values of the pilot, pulse, and noise, and the full estimator is detailed in Appendix C. We use the SNR estimate to compute $\zeta=\alpha_{n}\sqrt{n}=\left(\mathrm{SNR}\sqrt{1+r_{p/q}^{4}}\right)^{-1}\sqrt{32\log\left(\frac{1}{1-2\delta^{2}}\right)n_{s}}$. We set $\delta=0.05$ and use the result to compute $\alpha_{n}$ for each value of $T$. This is used to generate random vectors $\vec{t}$ with pulse locations used by Alice to transmit for each trial.
III-F Results
Fig. 9 plots Bob’s receiver’s performance vs. transmission duration $T$. Bob uses $\vec{t}$ to estimate bits only in the pulse locations that Alice uses for transmission, per sparse coding described in Section II-C. We report the decoding error probability $p_{e,\mathrm{bsc}}^{(b)}$ estimated by averaging over $N=150$ trials using the right ordinate of Fig. 9. We observe that $p_{e,\mathrm{bsc}}^{(b)}\approx 0.455$ throughout our experiments. Using the left ordinate we report the corresponding estimate of the total number of transmissible covert bits $B_{\rm bsc}(n)$ using the equality in (8). Fitting a line with slope of one-half to the log-log plot results in the coefficient of determination $R^{2}=0.91$, indicating the SRL-scaling that we expect.
Fig. 10 presents Willie’s detector performance. We employ the SNR estimation strategy described in Appendix C and plot using the right ordinate. We then obtain the lower bound on Willie’s probability of error $p_{e}^{(w)}$ by computing the upper bound on Hellinger distance between the probability distributions of his received signals under each hypothesis in Appendix A. We plot this estimate using the left ordinate and note that it is very conservative compared to the covertness requirement $p_{e}^{(w)}\geq 1-\delta=0.45$ as this is, effectively, a lower bound (11) on a lower bound (9). Because our initial SNR estimate closely matches the value measured during the experiments as $T$ increases, Alice can transmit while keeping the adversary’s error probability close to the desired level of $p_{e}^{(w)}=0.45$. This empirically demonstrates the viability of the SRL-governed sparse signaling approach.
@@ III-F Results
Figure 9: Bob’s receiver performance. Total number $B_{\mathrm{bsc}}(n)$ of reliably-decodable bits (left ordinate) and decoding error probability $p_{e,\mathrm{bsc}}^{(b)}$ (right ordinate) are plotted vs. transmission duration $T$, with 95% confidence intervals as error bars given $N=150$ trials per datapoint. Abscissa and left ordinate use a logarithmic scale; right ordinate uses a linear scale. Number of channel uses is $n=f_{s}T$.
Figure 10: Willie’s detector performance. Detection probability error $p_{e}^{(w)}$ (left ordinate) and SNR in dB (right ordinate) are plotted vs. transmission duration $T$, with 95% confidence intervals shown as error bars given $N=150$ trials per datapoint. Note that the left ordinate ranges from 0.40 to 0.475, and the right ranges from -20 dB to -50 dB. The abscissa is on a logarithmic scale with $\mathrm{SNR}\triangleq\frac{h_{w}^{2}c_{q}^{2}}{\sigma^{2}_{w}}$
IV Conclusion, Discussion, and Future Work
@@
We demonstrate the first implementation of SRL-based covert communication in RF domain. We employ SDRs, which are not specifically designed for covert communication. We have to address significant challenges:
@@
  • Dynamic range and ADC/DAC granularity limitations: The UBX-160 daughterboards use a 16-bit DAC for transmission and a 14-bit ADC for reception. While Alice can generate high-resolution low-power pulses, Bob’s lower ADC precision limits his ability to distinguish weak signals from noise. Ensuring detection of pulses after digitization leads to minimum transmission power, requiring the use of sparse coding. This, in turn, needs tight time synchronization, which we demonstrate.
@@
  • Time and frequency synchronization: The oven-controlled crystal oscillators (OCXOs) in USRP X310s lack the frequency accuracy and stability needed to decode sparsely transmitted covert symbols. Furthermore, Bob has to know precisely when the Alice begins transmitting. We utilize a common reference (OctoClock) and a non-covert preamble, however, in practice, GPS or a highly stable time source such as atomic clock can be used for both disciplining the local oscillators and timing information. We will employ these in follow-on experiments.
@@
  • Data volume: Our capture for $N=150$ trials amounts to approximately $3$ h of baseband data at a sampling rate of $f_{s}=12.5$ MHz, using 64-bit in-phase and quadrature samples for each receiver (Bob and Willie). This produces roughly $400$ GB of data. Although real-time processing is performed on the compute cluster, both computational and storage demands must be reduced for practical deployment.
  • Decoding performance degrades for longer durations: Transmissions with duration $T>4$ suffer from significant increase in Bob’s decoding error probability. Possible causes include LO instability and phase errors that our scheme does not correct. We will investigate means to mitigate this in future work.
@@
  • Continuous noise injection: Maintaining an always-on, spectrally flat artificial noise floor is limited by the maximum DAC sampling rate the system can sustain: sampling too fast can overwhelm the host CPU and lead to buffer under-flows. In the future, we will explore using more than one radio to emulate noise as well as more natural noise sources by experimenting “in the wild.”
@@
Indeed, our proof-of-concept experiment raises many theoretical and experimental research questions. In the short term, we plan to address some of them by improving our ORBIT-based design as follows:
@@
  • We will estimate Willie’s detection error probability directly by estimating the output distributions for the optimal test statistics, as is done in [20].
  • We will investigate the optimal constellation satisfying the Hellinger distance covertness criterion. There is evidence for QPSK being such constellation, as it is optimal for complex-valued circularly-symmetric AWGN under relative entropy criterion (by extension from the BPSK case for real-valued AWGN channels examined in [2, 9, 10]). We note that this is an active topic of investigation in quantum information theory [27].
Additionally, AWGN channel model provides only a zeroth-order approximation to practical RF channels. Thus, SRL-based covert communication needs to be validated in a more realistic, dynamic environment. We plan on evolving our system to be independent rather than centrally-controlled. This would allow not only experimentation “in the wild” but also exploration of covert networks. Furthermore, we plan on relaxing assumption on adversary’s capabilities and studying the impact of, e.g., lack of precise knowledge of the timing of the transmission and pulse shape used.

### v2

@@ III-C Pulse Optimization
To maximize the number of covert bits that can be reliably conveyed, we jointly tune four pulse-design parameters: the pilot-pulse length $n^{(p)}_{s}$, data-pulse length $n^{(q)}_{s}$, pilot-pulse magnitude $c_{p}$, and data-pulse magnitude $c_{q}$. We model the relationship between these parameters and the attainable covert bit count as a black-box objective, apply Bayesian optimization to efficiently navigate the parameter space, and use our hardware testbed to measure the covert bits delivered for each candidate configuration.
Figure 8: Data-packet structure (durations in samples). A 15 400-sample preamble is created by concatenating five 13-bit Barker sequences, BPSK-modulating them, zero-padding each to 200 samples, and convolving the result with a 2 401-sample root-raised-cosine filter. This is immediately followed by a one-second ($f_{s}$-sample) silence interval to measure the baseline noise level. The subsequent $T$-second, $T\times f_{s}$-sample segments encode the two parts of the experimental trial. In the Alice-on segment, each shaded pulse slot begins with a strictly positive, real-valued Gaussian pilot pulse (for phase reference), immediately followed by a Gaussian-shaped, QPSK-modulated data pulse carrying randomly generated information bits. The indices of the occupied slots are drawn uniformly at random and stored in $\vec{t}$ (see Section II-C). In the subsequent Alice-off segment, she remains silent.
Bayesian optimization iteratively explores the four-dimensional design space $\vec{d}=\left(n^{(p)}_{s},n^{(q)}_{s},c_{p},c_{q}\right)$ by maintaining a surrogate model of the unknown mapping from $\vec{d}$ to attainable covert bit count. We leverage the Ax library in Python [30] to perform this optimization. The three steps that follow are the standard Bayesian optimization loop [25, Ch. 31]; we restate them here.
@@
  • Measure : using $\vec{d}_{u}$, the radio transmits a modified version of the data packet (Section III-D): every fifth pulse slot in the Alice-on segment carries a random bit, and the Alice-off segment is removed. The number of attainable covert bits is recorded over $100$ one-second trials. The average number of attainable covert bits for a given parameter set is $r=2\,C_{\mathrm{bsc}}\,n_{p}\,\alpha_{n}$ bits over $n$ uses of the channel. Here, $n$ corresponds to the number of samples available in a one-second transmission. This yields a sample mean, $\bar{r}_{u}$, and a standard error, $s_{\bar{r}_{u}}$.
@@
After 200 iterations we use $\vec{d}^{\star}=\operatorname*{arg\,max}_{1\leq l\leq 200}\bar{r}_{l},$ the configuration with the highest observed mean throughput.
We use a Gaussian-process surrogate, which is commonly adopted in Bayesian optimization. We briefly review it here. For a set of $l$ different input parameter configurations $\vec{d}_{1:l}=[\vec{d}_{1},\vec{d}_{2},...\vec{d}_{l}]$, the model predicts $l$ corresponding covert throughput values sampled from the distribution $\mathcal{GP}\left(0,\,k(\vec{d}_{1:l},\vec{d}_{1:l})\right)$ [50], where $k$ is a covariance function, which is the Matérn-5/2 function [50, Ch. 4] here. This model can also be used to predict the throughput given priors, i.e., given u1u-1 tuples $(\vec{d},\bar{r},s_{\bar{r}})$, the model predicts the throughput $\hat{r}_{u}$ for some new parameter set $\vec{d}_{u}$. This prediction is distributed as $\hat{r}_{u}\mid\bar{r}_{1:(u-1)},\,\vec{d}_{1:u},\,s_{\bar{r}_{1:(u-1)}}\sim\mathcal{N}(\mu_{u},\sigma_{u}^{2})$, where $\mu_{u}$ and $\sigma_{u}$ can be directly calculated. At iteration $u$, configuration $d_{u}$ that maximizes expected improvement $\mathbb{E}[\hat{r}_{u}^{\prime}\mid\hat{r}_{u}^{\prime}>\bar{r}^{\star}_{1:(u-1)}]$ for a candidate parameter configuration $d_{u}^{\prime}$ is proposed for measurement, where $\hat{r}^{\star}_{1:(u-1)}$ is the highest measured throughput for the first u1u-1 trials, i.e., $r^{\star}_{1:(u-1)}=\max_{1\leq v\leq(u-1)}\bar{r}_{v}$.
The resulting optimum $(n^{(p)}_{s}=26,n^{(q)}_{s}=34,c_{p}=2.9765,c_{q}=3.521)$ is adopted for the experiments that follow.
@@ III-D Transmission Structure for Experimental Trials
In each experimental trial, Alice transmits the four-segment packet shown in Fig. 8: a non-covert preamble, a baseline-estimation gap, an Alice-on segment, and an Alice-off (noise-only) segment. An AWGN generator operates continuously throughout the experiment. The Alice-on and Alice-off segments last $T$ seconds each, while the preamble marks the start of each trial and facilitates finding Alice’s transmission in Bob’s and Willie’s traces.
The preamble consists of the 13-bit Barker sequence repeated five times (65 bits in total). Each bit is BPSK-modulated and occupies 200 samples; the first sample carries the modulated symbol, followed by 199 zero-padding samples. The resulting 13 000-sample sequence is then pulse-shaped with a root-raised-cosine (RRC) filter with roll-off factor $\beta=0.35$, spanning 12 symbols and represented digitally by $12\times 200+1=2401$ taps. After filtering, the preamble length is $65\times 200+(2401-1)=15,400$ samples. The baseline-estimation gap is one second ($f_{s}$-samples) long, and is used to monitor noise power level throughout the experiment.
@@ III-E Experiment design
We select eight logarithmically-spaced values of Alice’s transmission duration $T\in[10^{-0.3},10^{0.75}]$ s. For each value of $T$, we conduct $N=10^{3}$ independent trials using distinct transmission packets described in Section III-D. Before we begin our experiment, we estimate Willie’s SNR using a $20$ two-second calibration transmission with a modified version of the data packet from Section III-D: every fifth pulse slot in the Alice-on segment is used to transmit a random bit, and the Alice-off segment is deleted. The SNR is computed from the mean power values of the pilot, pulse, and noise, and the full estimator is detailed in Appendix F. We use the SNR estimate to compute $\zeta=\alpha_{n}\sqrt{n}=$, where $n=n_{p}n_{s}$ and $\alpha_{n}$ is in (13). We set $\delta=0.07$ and use the result to compute $\alpha_{n}$ for each value of $T$. This models Alice following the SRL. We also evaluate two scenarios when Alice carelessly does not follow the SRL, where $\alpha_{n}=\{6.6\times 10^{-4},3.31\times^{10^{-3}}\}$, the first corresponding to $\alpha_{n}$ for the SRL at $T=10^{-0.3}$ s. We use these three values of $\alpha_{n}$ to generate random vectors $\vec{t}$ with pulse locations used by Alice to transmit for each trial.
@@
Figure 9: Bob’s receiver performance. We plot the total number $B_{\mathrm{bsc}}(n)$ of reliably-decodable bits for $n=f_{s}T$ channels uses vs. transmission duration $T$ on a log-log scale, with 95% confidence intervals as error bars give $N=10^{3}$ trials per datapoint. $R^{2}$ is the coefficient of determination.
IV Results
IV-A Covert Throughput
Fig. 9 plots Bob’s receiver’s performance vs. transmission duration $T$. Bob uses $\vec{t}$ to estimate bits only at the pulse locations Alice uses for transmission, per the sparse coding described in Section II-C. We report the estimates of the total number of transmissible covert bits $B_{\rm bsc}(n)$ by averaging over $N=10^{3}$ trials using the equality in (8). Per Section III-E, Alice either follows the SRL or does not. We fit lines with slopes of either one-half or unity to the corresponding log-log plot results. We calculate the coefficient of determination $R^{2}$, which is the proportion of the variation in $B_{\rm bsc}(n)$ explained by the square-root or linear dependence on $T$ [20, Ch. 1.3]. Our values of $R^{2}>0.99$ indicate the scaling results that we anticipate. While careless Alice can transmit significantly more bits, this makes her detectable, as discussed next.
IV-B Detectability
We employ two methods to assess the detectability of Alice’s transmission: directly by using a constant false-alarm rate (CFAR) strategy, and by evaluating the probability of error $p_{e}{(w)}$ in (9). For both approaches, we use the optimal detector and the radiometer discussed in Section II-E. We consider a scenario where the noise power level is static, which favors Willie. We leave the extension to a dynamic environment, which could significantly favor Alice [40], for future work.
1 CFAR Strategy
Suppose Willie has a fixed false alarm budget $P_{\text{FA}}^{\star}$ and can collect detector data when Alice is absent. He then estimates a CFAR threshold on his test statistic. We implement this as follows: for a each value of $T$, and each detector in Section II-E, we find a threshold $\tau^{\star}$ on a test statistic $S$ such that the number of trials where $S>\tau^{\star}$ given $H_{0}$ (i.e., the false alarm events) is $P_{\text{FA}}^{\star}N$. Thus, $\tau^{\star}$ corresponds to a threshold that yields empirical false alarm probability $P_{\text{FA}}=P_{\text{FA}}^{\star}$. We then report the corresponding fraction of trials when $S\leq\tau^{\star}$ given $H_{1}$, which estimates the probability of missed detection $P_{\text{MD}}$. We note that our estimates of $P_{\text{MD}}$ and its 95% confidence intervals depend on $\tau^{\star}$. It, in turn, depends on the measured noise power level, which is random. However, tight control of our environment and a large number of trials used to calculate $\tau^{\star}$ render it nearly deterministic. Fig. 10 presents the results for $P_{\text{FA}}^{\star}=0.1$. We compare the detectors at the end of this section.
2 Evaluation of $p_{e}^{(w)}$
Fig. 11 presents our estimates of Willie’s probability of error $p_{e}^{(w)}$ in (9). First, we obtain the lower bound on $p_{e}^{(w)}$ when Alice follows the SRL by computing the upper bound on the Hellinger distance between the probability distributions of his received signals under each hypothesis in Appendix A. This is, effectively, a lower bound (12) on a lower bound (10). We estimate the SNR for each trial as described in Appendix F, calculate the bound, and plot the average over $N=10^{3}$ trials on both subfigures in Fig. 11. The correspondence to the covertness requirement $p_{e}^{(w)}\geq\frac{1}{2}-\delta=0.43$ indicates that our system operates as expected and that the noise power level and loss are stable throughout our experiments.
We also investigate the performance of the optimal detector and radiometer described in Section II-E. For each trial, we estimate the threshold $\frac{\sqrt{n_{p}}\Delta\mu}{\sigma_{0}+\sigma_{1}}$ for each detector using Willie’s observations as described in Appendix F and calculate the estimate for $p_{e,\text{det}}^{(w)}$ using (36). This corresponds to Willie implementing the detectors using information on the squared magnitude $a_{w}^{2}$ of the gain on his channel from Alice (which we assume he has), as well as measurements of the background noise $\sigma_{w}^{2}$. We note that the Berry-Esseen estimation errors $\epsilon_{n_{p},0}$ and $\epsilon_{n_{p},1}$ are negligible due to $n_{p}$ being large. We report the averages for each detector over $N=10^{3}$ trials.
3 Detector Comparison
Both strategies show that Willie’s probability of error remains constant when Alice follows the SRL. When Alice is careless, the optimal detector’s probability of error decays fairly rapidly with time $T$. On the other hand, the radiometer does not perform well in our experiments, even when $\alpha_{n}$ is a constant (corresponding to careless Alice). [^4] In a separate experiment we used a larger constant $\alpha_{n}$, and confirmed that radiometer detects careless Alice. Nevertheless, this demonstrates how Alice can exploit shortcomings in Willie’s detector design to improve her covert communication capabilities.
(a) Optimal Detector
(b) Radiometer
Figure 10: Comparison of missed detection probability $p_{\text{MD}}$ under CFAR $P_{\text{FA}}^{\star}=0.1$ vs. transmission duration $T$ (log scale) for LABEL:sub@fig:opt_perf_cfar the optimal detector and LABEL:sub@fig:rad_perf_cfar the radiometer. Results are averaged over $N=10^{3}$ trials with 95% confidence intervals shown.
(a) Optimal Detector
(b) Radiometer
Figure 11: Comparison of detection error probability $p_{e}^{(w)}$ versus transmission duration $T$ (log scale) for LABEL:sub@fig:opt_perf the optimal detector and LABEL:sub@fig:rad_perf the radiometer. Results are averaged over $N=10^{3}$ trials with 95% confidence intervals shown. The analytical lower bound from (12) is calculated using the power from the same trials, but its error bars are negligible and omitted for clarity.
V Conclusion, Discussion, and Future Work
@@
We demonstrate the first implementation of SRL-based covert communication in the RF domain. We employ SDRs, which are not specifically designed for covert communication. We have to address significant challenges:
@@
  • Dynamic range and ADC/DAC granularity limitations: The UBX-160 daughterboards use a 16-bit DAC for transmission and a 14-bit ADC for reception. While Alice can generate high-resolution low-power pulses, Bob’s lower ADC precision limits his ability to distinguish weak signals from noise. Ensuring pulse detection after digitization requires minimal transmission power, which necessitates sparse coding. This, in turn, needs tight time synchronization, which we demonstrate.
@@
  • Time and frequency synchronization: The oven-controlled crystal oscillators (OCXOs) in USRP X310s lack the frequency accuracy and stability needed to decode sparsely transmitted covert symbols. Furthermore, Bob has to know precisely when Alice begins transmitting. We use a common reference (OctoClock) and a non-covert preamble; however, in practice, GPS or a highly stable time source, such as an atomic clock, can be used for both disciplining the local oscillators and for timing information. We will employ these in follow-on experiments.
@@
  • Data volume: Our capture for $N=10^{3}$ trials amounts to approximately $72$ h of baseband data at a sampling rate of $f_{s}=12.5$ MHz, using 64-bit in-phase and quadrature samples for each receiver (Bob and Willie). This produces roughly $10$ TB of data. Although real-time processing is performed on the compute cluster, both computational and storage demands must be reduced for practical deployment.
@@
  • Continuous noise injection: Maintaining an always-on, spectrally flat artificial noise floor is limited by the maximum DAC sampling rate the system can sustain: sampling too fast can overwhelm the host CPU and lead to buffer under-flows. In the future, we will explore using multiple radios to emulate noise and more natural noise sources by experimenting “in the wild.”
@@
Indeed, our proof-of-concept experiment raises many theoretical and experimental research questions. In the short term, we plan to address some of them by improving our COSMOS-based design as follows:
@@
  • We will look for an optimal constellation satisfying the covertness criteria in (10) and (12). There is evidence for this being QPSK, as it is optimal for complex-valued circularly-symmetric AWGN under the criterion in (11) (by extending the BPSK case for real-valued AWGN channels in [10, 14, 47]). This is also actively investigated in quantum information theory [49].
Additionally, the AWGN channel model provides only a zeroth-order approximation to practical RF channels. Thus, SRL-based covert communication needs to be validated in a more realistic, dynamic environment. We plan on evolving our system to be independent rather than centrally controlled. This would allow not only experimentation “in the wild” but also exploration of covert networks. Furthermore, we plan to relax the assumption on the adversary’s capabilities and study the impact of, e.g., the lack of precise knowledge of the timing of transmission and the pulse shape used.